OpenStudy (anonymous):

Which equation has a graph parallel to the graph of 9x + 3y = –22

OpenStudy (anonymous):

parallel lines have the same slop :) so first we have to put the equation in slope intercept form :) Do you know how to do that ?

OpenStudy (anonymous):

sorta

OpenStudy (anonymous):

3y=-9x-22 y=-3x-(22/3)

OpenStudy (anonymous):

so all i had to do was isolate the y?

OpenStudy (anonymous):

Now, you know that if a line is gonna be parallel to the equation, it will have the same slope :)

OpenStudy (anonymous):

WAIT! the problem isnt finished ! LOL

OpenStudy (anonymous):

im sorry x)

OpenStudy (anonymous):

and the slope is 3 right?

OpenStudy (anonymous):

lollll its okay :P Soo...like i was saying...parallell lines have the same slope :) so whats the slop going to be ?

OpenStudy (anonymous):

NEGATIVE 3 :)

OpenStudy (anonymous):

sorry -3 :)

OpenStudy (anonymous):

its fine :) okay so does it give you point ?

OpenStudy (anonymous):

yeah :) thanks

OpenStudy (anonymous):

nooo lol i mean point like x and y like (x,y) because you need points to solve lol

OpenStudy (anonymous):

sorry im confusing >.<

OpenStudy (anonymous):

to find the points dont i replace x with 0 to get y and replace y with 0 to get x

OpenStudy (anonymous):

okay try that but usually problems like that come with points lol

OpenStudy (anonymous):

this one didnt x)

OpenStudy (anonymous):

thank you for stating the obvious ;) LOL

OpenStudy (anonymous):

lol i got y=-7.3 and x=-2.4

OpenStudy (anonymous):

okay so now you use point slope form which is y-y1=m(x-x1) sounds familiar ?

OpenStudy (anonymous):

yeah :)

OpenStudy (anonymous):

okay so now just plug everything in :)

OpenStudy (anonymous):

y+7.3=-3(x+2.4)

OpenStudy (anonymous):

good now make it slope intercept form :) y+7.3 = -3x -7.2 y=-3x -14.5 :)

OpenStudy (anonymous):

:) thanks i got the same

OpenStudy (anonymous):

yay :) good jobbbb

OpenStudy (anonymous):

:)